OpenStudy (anonymous):

Amistre64, are you out there today?

OpenStudy (amistre64):

im lucy in the sky..with diamonds :)

OpenStudy (amistre64):

koo-koo-kachoo

OpenStudy (anonymous):

I am so relieved! Need your help, if you don't mind giving it again today.

OpenStudy (amistre64):

i dunno, i was kinda busy watching the butterflies fight over a flower in the backyard..

OpenStudy (amistre64):

of course i picked all the other flowers just to instigate the fight, but they dont know that lol

OpenStudy (anonymous):

If all I had to do was watch butterflies!...Let me try to insert the equation here. Just give me a sec...

OpenStudy (anonymous):

\[\sqrt{144-(6\sqrt{2}})^{2}\]

OpenStudy (amistre64):

=0

OpenStudy (anonymous):

Can you help me figure out how to get the (6sqrt2*2) figured out, please.

OpenStudy (anonymous):

What are the rules for that?

OpenStudy (amistre64):

(ab)^2 = a^2 b^2 (6.sqrt(2))^2 = 6^2 [sqrt(2)]^2 36.2 = 72 sqrt(144-72) = sqrt(72)

OpenStudy (amistre64):

sqrt(abc) = sqrt(a) sqrt(b) sqrt(c) 9*4*2 = 72 sqrt(9) sqrt(4) sqrt(2) = 3.2 sqrt(2) = 6sqrt(2)

OpenStudy (anonymous):

Ok, I am trying to figure that out, how you got it, on my paper. But meanwhile the whole problem is 2(-6sqrt2)*(what i asked you to help me figure out...the sqrt of 144-(6sqrt2)^2. got that or should I try to insert the equation?

OpenStudy (amistre64):

your keeping secrets from me :) post the whole problem so that i can see the whole picture and come to a complete solution; instead of trying to fit into it half way thu :)

OpenStudy (anonymous):

gotcha! Give me a sec here...

OpenStudy (anonymous):

\[2(-6\sqrt{2)}\sqrt{144-(6\sqrt{2}})^{2}\]

OpenStudy (anonymous):

no secrets! promise!

OpenStudy (amistre64):

ok... i showed the results for the last factor ;) so yes, we continue with the muliplications. the innerworkings to remember here tho is that radicals act like fancy looking variables when we add; but act like real numbers when we multiply...

OpenStudy (amistre64):

2.-6sqrt(2).6sqrt(2) ; rearrange to show steps 2.-6.6.sqrt(2).sqrt(2) ; multiply the integers to clean up the scene -72.sqrt(2).sqrt(2) ; since sqrt(a).sqrt(b) = sqrt(ab) we continue thusly -72.sqrt(2.2) ; multiply the 2.2 -72.sqrt(4) ; look up inthe table of sqrts in the back of the book to determine what the sqrt(4) is........ sqrt(4) = 2; insert to in place of sqrt(4) -72.2 ; multiply using an abacus... -1(140 + 4) ; add the innards -1(144) ; and squeeze the -1 inside to get........ -144 TADA!!

OpenStudy (anonymous):

TADA my bootie...tada, oh pls. If I only were 1/16 as smart as you are! Again, I GET IT! Thank you so much for your time! Are you a math teacher or something!?

OpenStudy (amistre64):

im an 'or something' :)

OpenStudy (anonymous):

but you use words like "thusly"; part math guru/part english teacher?

OpenStudy (amistre64):

im old, so i have to start using old english ;)

OpenStudy (anonymous):

I would guess, by your user name, that you are 46 or 47. But you don't have to answer. I am 41, and this is all greek to me. Age does not help the brain, maturely speaking, figure things like this out. If you don't know it, you just don't know it and are in loads of trouble!

OpenStudy (amistre64):

im the class of 1990... borned in the backwoods of lakeland florida in the year 1972 :)

OpenStudy (anonymous):

So you're younger than me and you say you're old!? shame on you! I wish I had your brains. I'm sure I will be back here today sometime to pick your brain if you will be around. Thanks! Again and again, thanks, thanks, thanks! Have a great morning in Florida!

OpenStudy (amistre64):

youre welcome :) enjoy ;)

OpenStudy (anonymous):

Enjoy Calculus? Ha! Actually, when everything comes together, I DO enjoy it! In fact, I love to see the puzzle pieces all fall into the right spot! That's a thrill to me! Boy, i am old! ; )

OpenStudy (amistre64):

an exacto knife helps to when you know your right lol. I thought it said 1000 pieces!!

OpenStudy (anonymous):

Tell me about it! Ttysoon!